solve for the unknown:
2/3 /10 = 1.8/ r
r= ?
These are two sets of fractions with an = sign in the middle.
I really don't know where to start. Maybe cross multiply?

Answer by checkley77(12844)

(2/3)/10=1.8/r
2/3r=1.8*10
2r/3=18
2r=3*18
2r=54
r=54/2
r=27 ans.
Proof:
(2/3)/10=1.8/27
2/3*27=1.8*10
18=18